# Local Causal Discovery for Structural Evidence of Direct Discrimination

### v1

@@ Appendix D Empirical Evaluation Methods
    LDECC: With post-processing, we retain the variables labeled as parents of the target unioned with the variables that are unoriented.
@@
    LDP: Without post-processing, we retain $\mathbf{Z}_{1}\cup\mathbf{Z}_{\textsc{POST}}\cup\mathbf{Z}_{4}$. As our experimental setting requires that there are no descendants of the outcome, the superset $\mathbf{Z}_{\textsc{POST}}\coloneqq\mathbf{Z}_{2\in de(Y)}\cup\mathbf{Z}_{3}\cup\mathbf{Z}_{6}$ defined in Maasch et al. [48] will now only contain $\mathbf{Z}_{3}$.
@@
    LDPpa: We post-process the LDP output to only retain the variables that are not conditionally independent of $Y$ given the remaining members of the set.
Appendix E Empirical Results
@@
Figure E.1: Total runtime and total conditional independence tests performed with respect to total edges and total nodes, using an oracle on random directed acyclic Erdős-Rényi graphs [23].
Figure E.2: Mean runtime for the Asia [45] and Sachs [74] benchmarks from bnlearn (https://www.bnlearn.com/bnrepository/) [76] ($\chi^{2}$ tests; $\alpha=0.001$; 95% confidence intervals in shaded regions). For Asia (either $\to$ dysp), $|\mathbf{Z}|=6$ and $|\mathbf{A}_{\text{{DE}}}|=1$. For Sachs (erk $\to$ akt), $|\mathbf{Z}|=11$ and $|\mathbf{A}_{\text{{DE}}}|=1$. Raw data are reported in Tables E.2 and E.4.
@@
Figure E.3: Mean weighted CDE estimates over 100 replicates per sample size for linear-Gaussian instantiations of Figure D.1 (95% confidence intervals in shaded regions). Ground truth weighted CDE is denoted by the dashed line. LD3 used parametric Fisher-z tests ($\alpha=0.01$), as the data generating process was known for this simulation. Estimates were obtained with linear regression (https://scikit-learn.org/). Note: This pipeline is for ease of explication in the illustrative setting only. For unknown data generating processes, we do not recommend this pipeline due to restrictive parametric assumptions.
@@
Figure E.4: SDC accuracy, $\mathbf{A}_{\text{{DE}}}$ F1, precision, and recall across sample sizes when SDC = 0 and SDC = 1. Values were averaged over 100 replicates of linear-Gaussian instantiations of Figure D.1, with 95% confidence intervals in shaded regions. LD3 used parametric Fisher-z tests ($\alpha=0.01$), as the data generating process was known for this simulation.
@@
Table E.1: Empirical performance on the Asia (Lung Cancer) Bayesian network dataset [44] from bnlearn [76]. Binary data were evaluated with $\chi^{2}$ independence tests ($\alpha=0.001$; 95% confidence intervals in brackets). Unexpected decreases in baseline performance for increasing sample size might be due to the difficulty of this dataset for structure learning, as noted in the bnlearn documentation: “Standard learning algorithms are not able to recover the true structure of the network because of the presence of a node (E) with conditional probabilities equal to both 0 and 1" (https://www.bnlearn.com/documentation/man/asia.html).
@@
Table E.2: Total number of independence tests performaed and runtimes in seconds (95% confidence intervals in brackets).
@@
Table E.3: Performance on the Sachs protein dataset [74] from bnlearn [76]. Discrete data were evaluated with $\chi^{2}$ independence tests ($\alpha=0.001$; 95% confidence intervals in brackets).
@@
Table E.4: Total number of independence tests and runtimes (95% confidence intervals in brackets).
Table E.5: SDC assessment and weighted CDE estimation with double machine learning using results from PC Algorithm [84], as implemented by Kalisch & Buhlmann [38]. We define $\mathbf{A}_{\text{{DE}}}$ using the union method (PC; Appendix D), though results for $\mathbf{A}_{\text{{DE}}}$ were exactly the same using PC.
Unos Policy (2017–2019) Unos Policy (2020–2022) LDECC ($\alpha=0.01$) LDECC ($\alpha=0.05$) LDECC ($\alpha=0.01$) LDECC ($\alpha=0.05$) $\mathbf{A}_{\text{{DE}}}$ payment method payment method, region active exception case, diagnosis, initial MELD, initial age, recipient blood type diagnosis, recipient blood type, region SDC 1 1 1 1 WCDE 0.049 [0.032,0.065] 0.048 [0.032,0.064] 0.032 [0.019,0.046] 0.032 [0.017,0.048] $p$-value 0.000 0.000 0.000 0.000
Table E.6: SDC assessment and weighted CDE estimation with double machine learning using results from LDECC using the union method for selection $\mathbf{A}_{\text{{DE}}}$ (LDECC; Appendix D) [31]. Results for $\mathbf{A}_{\text{{DE}}}$ were almost concordant using LDECC, with the exception of 2020-2022 when $\alpha=0.05$ (where the intersection method excluded diagnosis).
Table E.7: Total number of independence tests performed and runtimes in seconds for LD3 vs PC [38] and LDECC [31] on the liver transplant allocation dataset. While PC is worst-case exponential in time complexity [84], LDECC is worst-case exponential for some structures and polynomial for others [31]. This could explain the wide variation in LDECC runtimes across experimental settings.
Appendix F Liver Fairness

### v3

@@ D.2 Baseline Evaluation
    PC: Retain the union of parent sets across DAGs in the MEC.
@@
    MB-by-MB: Retain the variables labeled as parents of the target.
@@
    LDECC: Retain the variables labeled as parents of the target.
  7. 7.
    LDECC: Retain the variables labeled as parents of the target unioned with the variables adjacent to $Y$ that are unoriented in the MEC.
  8. 8.
    LDP: Retain $\mathbf{Z}_{1}\cup\mathbf{Z}_{\textsc{POST}}\cup\mathbf{Z}_{4}$. As our experimental setting requires that there are no descendants of the outcome, the superset $\mathbf{Z}_{\textsc{POST}}\coloneqq\mathbf{Z}_{2\in de(Y)}\cup\mathbf{Z}_{3}\cup\mathbf{Z}_{6}$ defined in Maasch et al. (2024) will now only contain $\mathbf{Z}_{3}$.
  9. 9.
@@
    LDPpa: Retain the variables that are not conditionally independent of $Y$ given the remaining members of the set.
For the Sangiovese benchmark, we compared LD3 to NOTEARS, DirectLiNGAM, MB-by-MB, LDECC, PC, and LDPpa. For Asia and Sachs, we compared LD3 to PC, PC, LDECC, LDECC, LDP with no post-processing, and LDPpa to emphasize the impacts of different post-processing methods.
D.3 Erdős-Rényi Graphs
@@
Figure D.2: Total runtime and total CI tests performed with respect to total edges and total nodes, using an oracle on random directed acyclic Erdős-Rényi graphs (Erdős, Rényi et al. 1960).
D.4 Weighted CDE Estimation
Results
To evaluate SDC and WCDE correctness, we tested LD3 on linear-Gaussian instantiations of an 18-node DAG with an M-structure (i.e., a member of $\mathbf{Z}_{2\notin de(Y)}$; Ding and Miratrix 2015) and multiple members of $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{4}$ (Figure D.1). The WCDE was estimated using linear regression with $\mathbf{A}_{\text{{DE}}}$ as covariates. $\mathbf{A}_{\text{{DE}}}$ F1, precision, and recall and SDC accuracy converged toward 100% as sample size increased. The WCDE converged toward the true direct effect with low variance (Figures D.3, D.4).
@@ Results
Figure D.3: Mean WCDE estimates over 100 replicates per sample size for linear-Gaussian instantiations of Figure D.1 (95% confidence intervals in shaded regions). Ground truth WCDE is denoted by the dashed line. LD3 used parametric Fisher-z tests ($\alpha=0.01$), as the data generating process was known for this simulation. Estimates were obtained with linear regression (https://scikit-learn.org/). Note: This pipeline is for ease of explication in the illustrative setting only. For unknown data generating processes, we do not recommend this pipeline due to restrictive parametric assumptions.
@@
Figure D.4: SDC accuracy, $\mathbf{A}_{\text{{DE}}}$ F1, precision, and recall across sample sizes when SDC = 0 and SDC = 1. Values were averaged over 100 replicates of linear-Gaussian instantiations of Figure D.1, with 95% confidence intervals in shaded regions. LD3 used parametric Fisher-z tests ($\alpha=0.01$), as the data generating process was known for this simulation.
D.5 Unnecessary Adjustment: Estimator Variance and VAS Interpretability
[…]
Figure D.5: The true $\mathbf{A}_{\text{{DE}}}$ for $X$ and $Y$ is in gray.
$\mathbf{A}_{\text{{DE}}}$ All $\mathbf{Z}$
$n$ Mean Variance Mean Variance
100 6.98 0.83 6.97 7.79
1000 7.17 0.86 7.28 8.27
10000 6.91 1.16 7.25 9.04
Table D.2: Given the DAG in Figure D.5, adjusting for $\mathbf{Z}$ inflates WCDE variance relative to $\mathbf{A}_{\text{{DE}}}$ in finite samples (size $n$). True WCDE was 7. Data generating process was linear-Gaussian. WCDE estimation used linear regression with 100 replicates per sample size.
All $\mathbf{Z}$ True $\mathbf{A}_{\text{{DE}}}$ Pred $\mathbf{A}_{\text{{DE}}}$
$n$ Mean Variance Mean Variance Mean Variance $\mathbf{A}_{\text{{DE}}}$ F1
500 0.239 0.052 0.347 0.004 0.344 0.004 0.99 [0.98,1.0]
1000 -0.011 0.038 0.35 0.003 0.349 0.003 0.99 [0.98,1.0]
10000 0.151 0.013 0.345 0.000 0.344 0.000 0.99 [0.98,1.0]
Table D.3: Statistical Efficiency. WCDE estimate mean and variance when adjusting for all $\mathbf{Z}$, the true $\mathbf{A}_{\text{{DE}}}$ defined by our graphical criterion (True $\mathbf{A}_{\text{{DE}}}$), and the predicted $\mathbf{A}_{\text{{DE}}}$ returned by LD3 (Pred $\mathbf{A}_{\text{{DE}}}$) for the DAG in Figure D.5. Variance using all $\mathbf{Z}$ was at least 12.6×12.6$\times$12.6 × higher than using Pred $\mathbf{A}_{\text{{DE}}}$. F1 was high and variance for Pred $\mathbf{A}_{\text{{DE}}}$ was equal to the variance using the ground truth $\mathbf{A}_{\text{{DE}}}$. Variables were categorical with quadratic causal functions. WCDE estimation used double machine learning using econml (https://econml.azurewebsites.net/). Means are over 100 replicates per sample size (95% confidence intervals in brackets). Though adjusting for all $\mathbf{Z}$ will not theoretically induce bias in this setting, we see that mean WCDE values do diverge (and sometimes flip sign).
$n$ $|\mathbf{Z}|$ $|\mathbf{A}_{\text{{DE}}}|$ Acc F1 Rec Prec Time Ecoli $1k$ 44 2.0 [2.0,2.0] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 0.016 [0.01,0.021] Andes $50k$ 221 4.5 [4.06,4.94] 1.00 [1.00,1.00] 0.95 [0.9,0.99] 0.91 [0.83,0.99] 1.00 [1.00,1.00] 0.76 [0.737,0.784]
Table D.4: Interpretability. $\mathbf{A}_{\text{{DE}}}$ discovered by LD3 can provide greater interpretability than adjusting for all $\mathbf{Z}$, especially in large DAGs. We used the binary Andes (Conati et al. 1997) and Gaussian Ecoli (Schäfer and Strimmer 2005) benchmarks from bnlearn (Scutari 2010). When the true parent set is small but total number of nodes is moderate (e.g., Ecoli) to large (e.g., Andes), forgoing structure learning by adjusting for $\mathbf{Z}$ is a missed opportunity to provide mechanistic insights (e.g., causal fairness mechanisms). In the fairness setting, we can draw direct discrimination conclusions about the members of $\mathbf{A}_{\text{{DE}}}$, but not $\mathbf{Z}$ alone. On average, eliminating extraneous variables took less than one second. Ecoli used Fisher-z tests ($\alpha=0.001$) and Andes used $\chi^{2}$ ($\alpha=0.01$). Values were averaged over ten replicates (95% confidence intervals in brackets).
D.6 bnlearn Benchmarks
$n$ Method Time CI tests Accuracy F1 Precision Recall 250 LD3 0.005 [0.003,0.008] 60.8 [60.41,61.19] 0.74 [0.71,0.78] 0.7 [0.65,0.76] 1.00 [1.00,1.00] 0.55 [0.48,0.62] LDECC 0.092 [0.063,0.12] 316.6 [220.98,412.22] 0.61 [0.55,0.67] 0.47 [0.33,0.6] 0.9 [0.7,1.0] 0.32 [0.2,0.43] MB-by-MB 0.092 [0.047,0.136] 268.4 [179.54,357.26] 0.63 [0.57,0.68] 0.5 [0.37,0.62] 0.9 [0.7,1.0] 0.35 [0.25,0.45] LDP 0.009 [0.006,0.012] 112.1 [90.82,133.38] 0.76 [0.72,0.81] 0.76 [0.71,0.8] 0.94 [0.87,1.0] 0.64 [0.59,0.68] PC 0.118 [0.099,0.136] 557.9 [524.73,591.07] 0.64 [0.58,0.69] 0.51 [0.39,0.64] 0.9 [0.7,1.0] 0.36 [0.27,0.46] DirectLiNGAM 0.142 [0.138,0.145] - 0.63 [0.55,0.7] 0.56 [0.43,0.69] 0.85 [0.75,0.95] 0.48 [0.32,0.63] NOTEARS 1.216 [0.985,1.447] - 0.79 [0.75,0.82] 0.79 [0.75,0.82] 0.91 [0.86,0.96] 0.7 [0.65,0.75] 500 LD3 0.006 [0.004,0.008] 60.8 [60.41,61.19] 0.85 [0.79,0.91] 0.84 [0.77,0.91] 1.00 [1.00,1.00] 0.74 [0.64,0.84] LDECC 0.137 [0.104,0.169] 458.2 [354.85,561.55] 0.68 [0.63,0.73] 0.6 [0.52,0.69] 0.98 [0.94,1.0] 0.45 [0.37,0.53] MB-by-MB 0.127 [0.113,0.141] 370 [328.41,411.59] 0.64 [0.61,0.68] 0.55 [0.48,0.61] 0.98 [0.94,1.0] 0.39 [0.32,0.46] LDP 0.01 [0.005,0.016] 84.3 [66.81,101.79] 0.78 [0.73,0.82] 0.77 [0.72,0.82] 0.94 [0.88,1.0] 0.66 [0.6,0.73] PC 0.181 [0.162,0.201] 879.6 [843.24,915.96] 0.61 [0.56,0.67] 0.47 [0.38,0.57] 1.00 [1.00,1.00] 0.32 [0.23,0.42] DirectLiNGAM 0.152 [0.144,0.16] - 0.56 [0.46,0.67] 0.42 [0.18,0.65] 0.43 [0.19,0.67] 0.41 [0.18,0.65] NOTEARS 1.092 [1.025,1.158] - 0.79 [0.76,0.82] 0.78 [0.75,0.81] 0.95 [0.9,1.0] 0.66 [0.63,0.7] 1000 LD3 0.009 [0.007,0.011] 61 [61,61] 0.97 [0.95,0.99] 0.99 [0.95,0.99] 1.00 [1.00,1.00] 0.95 [0.91,0.99] LDECC 0.217 [0.185,0.248] 669.3 [560.92,777.68] 0.66 [0.61,0.72] 0.57 [0.45,0.68] 1.00 [1.00,1.00] 0.41 [0.31,0.52] MB-by-MB 0.201 [0.175,0.228] 592.7 [538.57,646.83] 0.69 [0.66,0.72] 0.63 [0.58,0.68] 1.00 [1.00,1.00] 0.46 [0.41,0.51] LDP 0.006 [0.003,0.009] 75 [63.46,86.54] 0.95 [0.91,0.99] 0.95 [0.91,0.99] 0.99 [0.96,1.0] 0.92 [0.87,0.98] PC 0.289 [0.272,0.306] 1333.9 [1249.08,1418.72] 0.55 [0.5,0.6] 0.33 [0.21,0.45] 0.8 [0.54,1.06] 0.21 [0.13,0.29] DirectLiNGAM 0.366 [0.326,0.407] - 0.62 [0.55,0.69] 0.56 [0.4,0.73] 0.66 [0.51,0.8] 0.54 [0.34,0.73] NOTEARS 1.694 [1.571,1.818] - 0.8 [0.77,0.83] 0.79 [0.75,0.82] 1.00 [1.00,1.00] 0.65 [0.6,0.7]
Table D.5: Results on the Sangiovese benchmark from bnlearn (Scutari 2010), a linear-Gaussian model of Tuscan Sangiovese grape production (Magrini, Di Blasi, and Stefanini 2017). Ten replicate datasets were sampled at $n=[250,500,1000]$. Means are given with 95% confidence intervals in brackets. Exposure was pH and outcome was GrapeW. All constraint-based methods used Fisher-z tests ($\alpha=0.01$). NOTEARS hyperparameters were $\lambda_{1}=0.001$ and loss type $l_{2}$. DirectLiNGAM: https://lingam.readthedocs.io/; PC, LDECC, MB-by-MB: https://github.com/acmi-lab/local-causal-discovery; NOTEARS: https://github.com/xunzheng/notears; LDP: https://github.com/jmaasch/ldp.
Figure D.6: Total number of CI tests performed for the Asia (Lauritzen and Spiegelhalter 1988) and Sachs (Sachs et al. 2005) benchmarks from bnlearn (https://www.bnlearn.com/bnrepository/) (Scutari 2010) ($\chi^{2}$ tests; $\alpha=0.001$; 95% confidence intervals in shaded regions). For Asia (either $\to$ dysp), $|\mathbf{Z}|=6$ and $|\mathbf{A}_{\text{{DE}}}|=1$. For Sachs (erk $\to$ akt), $|\mathbf{Z}|=11$ and $|\mathbf{A}_{\text{{DE}}}|=1$. Raw data are reported in Tables D.7 and D.9.
@@ D.5 Unnecessary Adjustment: Estimator Variance and VAS Interpretability
Table D.6: Empirical performance on the Asia (Lung Cancer) Bayesian network dataset (Lauritzen and Spiegelhalter 1988) from bnlearn (Scutari 2010). Binary data were evaluated with $\chi^{2}$ independence tests ($\alpha=0.001$; 95% confidence intervals in brackets). Unexpected decreases in baseline performance for increasing sample size might be due to the difficulty of this dataset for structure learning, as noted in the bnlearn documentation: “Standard learning algorithms are not able to recover the true structure of the network because of the presence of a node (E) with conditional probabilities equal to both 0 and 1” (https://www.bnlearn.com/documentation/man/asia.html).
@@
Table D.7: Total number of independence tests performed and runtimes in seconds (95% confidence intervals in brackets).
@@
Table D.8: Performance on the Sachs protein dataset (Sachs et al. 2005) from bnlearn (Scutari 2010). Discrete data were evaluated with $\chi^{2}$ independence tests ($\alpha=0.001$; 95% confidence intervals in brackets).
@@
Table D.9: Total number of independence tests and runtimes (95% confidence intervals in brackets).
Appendix E Real-World Causal Fairness Experiments
Computing Resources
All causal fairness experiments used an Apple MacBook (M2 Pro Chip; 12 CPU cores; 16G memory).
E.1 COMPAS Recidivism Prediction
Data Preprocessing
We used the data file compas-scores-two-years.csv from ProPublica, [^14] https://github.com/propublica/compas-analysis retaining the following columns for our analyses:
  1. 1.
    race (float64)
  2. 2.
    sex (float64)
  3. 3.
    age_cat (float64)
  4. 4.
    juv_fel_count (int64)
  5. 5.
    juv_misd_count (int64)
  6. 6.
    juv_other_count (int64)
  7. 7.
    priors_count (int64)
  8. 8.
    c_charge_degree (float64)
  9. 9.
    two_year_recid (int64)
  10. 10.
    decile_score (int64)
  11. 11.
    race_binary (int64)
Race was filtered to retain African American ($n=3696$) and white individuals ($n=2454$). For each exposure-outcome pair, $|\mathbf{Z}|=7$. When exposure was race and outcome was decile_score, $\mathbf{Z}$ contained all features other than two_year_recid (as this temporally proceeds decile_score, violating A1). When exposure was race and outcome was two_year_recid, $\mathbf{Z}$ contained all features other than decile_score.
White ($n$ = 2454) African American ($n$ = 3696) Age 37.73 (12.76) 32.74 (10.86) Charge degree (felony) 0.6 (0.49) 0.69 (0.46) Charge degree (misdemeanor) 0.4 (0.49) 0.31 (0.46) General recidivism decile score 3.74 (2.6) 5.37 (2.83) Juvenile felony count 0.03 (0.3) 0.1 (0.51) Juvenile misdemeanor count 0.04 (0.28) 0.14 (0.61) Juvenile other count 0.09 (0.45) 0.14 (0.56) Priors count 2.59 (3.8) 4.44 (5.58) Sex (female) 0.23 (0.42) 0.18 (0.38) Sex (male) 0.77 (0.42) 0.82 (0.38) Two-year recidivism (no) 0.61 (0.49) 0.49 (0.5) Two-year recidivism (yes) 0.39 (0.49) 0.51 (0.5)
Table E.1: Summary statistics for African American versus white individuals in the COMPAS dataset after preprocessing. We report mean values (standard deviations).
Refer to caption Age Refer to caption Sex Refer to caption Charge Degree Refer to caption Priors Count Refer to caption Two-year Recidivism Refer to caption General Recidivism Decile Score
Figure E.1: Distributions for African American versus white individuals in the COMPAS dataset after preprocessing.
Figure E.2: Average number of CI tests performed and average time (seconds) per method on COMPAS experiments.
$DS$$CD$$R$$A$$JF$$PC$ (A) $\alpha=0.005$ $SDC=1$ $WCDE=0.694\;[0.548,0.839]$ $p=0.000$
$DS$$CD$$R$$A$$JF$$PC$ (B) $\alpha=0.01$ $SDC=1$ $WCDE=0.695\;[0.55,0.84]$ $p=0.000$
$DS$$CD$$R$$A$$JF$$PC$$S$$JM$ (C) $\alpha=0.05$ $SDC=1$ $WCDE=0.657\;[0.51,0.804]$ $p=0.000$
$TY$$CD$$A$$PC$$S$ (D) $\alpha=0.005$ $SDC=0$ $WCDE=0.003\;[-0.028,0.033]$ $p=0.87$
$TY$$CD$$A$$PC$$S$$R$$JM$ (E) $\alpha=0.01$ $SDC=1$ $WCDE=0.007\;[-0.024,0.037]$ $p=0.675$
$TY$$CD$$A$$PC$$S$$R$$JF$$JM$ (F) $\alpha=0.05$ $SDC=1$ $WCDE=0.005\;[-0.026,0.035]$ $p=0.766$
Figure E.3: LD3 results on the COMPAS dataset at different significance levels ($\alpha=[0.005,0.01,0.05]$). DAGs represent the predicted parent set for the outcome only, with all other relations abstracted away. (A–C) Exposure race (R; red) and outcome general recidivism decile score (DS; blue). Known parents of outcome DS are in yellow. (D–F) Exposure race and outcome two-year recidivism (TY; blue). A = age; CD = charge degree; DS = decile score; JF = juvenile felony count; JM = juvenile misdemeanor count; PC = priors count; R = race; S = sex; TY = two-year recidivism. Weighted CDE ($WCDE$) is reported with $p$-value ($p$) and 95% confidence intervals in brackets.
$DS$$A$$PC$$R$ (A) $\alpha=0.005$ $SDC=1$ $WCDE=0.709\;[0.564,0.853]$ $p=0.000$
$DS$$A$$PC$$R$ (B) $\alpha=0.01$ $SDC=1$ $WCDE=0.71\;[0.565,0.855]$ $p=0.000$
$DS$$A$$PC$$R$ (C) $\alpha=0.05$ $SDC=1$ $WCDE=0.71\;[0.565,0.855]$ $p=0.000$
$TY$$PC$$R$ (D) $\alpha=0.005$ $SDC=1$ $WCDE=0.058\;[0.029,0.088]$ $p=0.000$
$TY$$PC$$R$ (E) $\alpha=0.01$ $SDC=1$ $WCDE=0.059\;[0.029,0.089]$ $p=0.000$
$TY$$PC$$R$$A$$S$ (F) $\alpha=0.05$ $SDC=1$ $WCDE=0.004\;[-0.026,0.034]$ $p=0.797$
Figure E.4: PC results on the COMPAS dataset at different significance levels ($\alpha=[0.005,0.01,0.05]$). DAGs represent the predicted parent set for the outcome only, with all other relations abstracted away. (A–C) Exposure race (R; red) and outcome general recidivism decile score (DS; blue). Known parents of outcome DS are in yellow. (D–F) Exposure race and outcome two-year recidivism (TY; blue). A = age; DS = decile score; PC = priors count; R = race; S = sex; TY = two-year recidivism. Weighted CDE ($WCDE$) is reported with $p$-value ($p$) and 95% confidence intervals in brackets.
$DS$$A$$PC$$R$ (A) $\alpha=0.005$ $SDC=1$ $WCDE=0.711\;[0.566,0.856]$ $p=0.000$
$DS$$A$$PC$$R$ (B) $\alpha=0.01$ $SDC=1$ $WCDE=0.709\;[0.564,0.853]$ $p=0.000$
$DS$$CD$$R$ (C) $\alpha=0.05$ $SDC=1$ $WCDE=1.559\;[1.396,1.723]$ $p=0.000$
$TY$$A$$PC$ (D) $\alpha=0.005$ $SDC=0$ $WCDE=0.012\;[-0.018,0.042]$ $p=0.445$
$TY$$PC$$R$$A$$S$ (E) $\alpha=0.01$ $SDC=1$ $WCDE=0.006\;[-0.024,0.036]$ $p=0.693$
$TY$$A$$S$$R$ (F) $\alpha=0.05$ $SDC=1$ $WCDE=0.082\;[0.052,0.113]$ $p=0.000$
Figure E.5: LDECC results on the COMPAS dataset at different significance levels ($\alpha=[0.005,0.01,0.05]$). DAGs represent the predicted parent set for the outcome only, with all other relations abstracted away. (A–C) Exposure race (R; red) and outcome general recidivism decile score (DS; blue). Known parents of outcome DS are in yellow. (D–F) Exposure race and outcome two-year recidivism (TY; blue). A = age; CD = charge degree; DS = decile score; PC = priors count; R = race; S = sex; TY = two-year recidivism. Weighted CDE ($WCDE$) is reported with $p$-value ($p$) and 95% confidence intervals in brackets.
E.2 Liver Transplant Allocation